OpenStudy (anonymous):

Using complete sentences, explain how to find the equation of the line, in standard form and slope–intercept form, passing through (–1, 5) and (1, 9). Also compare the benefits of writing an equation in standard form to the benefits of writing an equation in slope–intercept form.

jimthompson5910 (jim_thompson5910):

What is the slope of the line through those two points?

OpenStudy (anonymous):

Slope of the line = 2

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so m = 2

jimthompson5910 (jim_thompson5910):

now turn to y = mx + b

jimthompson5910 (jim_thompson5910):

plug in m = 2 also plug in either point for x and y then solve for b

OpenStudy (anonymous):

Standard form will give you the intercepts of both x and y. For example, in 5x+4y=10, the x intercept would be 5 and the y intercept would be 4. As for the first part, find the slope using (y2-y1)/(x2-x1). Then using the slope to find the equation of the line using point slope form. Which would be y-y1=m(x-x1) using either of your points. So to solve your problem, using the points (-1,5) and (1,9) (9-5=4)/(1--1=2) slope = 2. Then we can use either -1,5 or 1,9 to plug into the point slope equation. y-9-2(x-1) y-9=2x-2 +9 to both sides y=2x+7

OpenStudy (anonymous):

Thank you both soooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooo much! ^^